Find the four roots of z⁴=-1


1
Expert's answer
2022-06-15T05:49:22-0400

"z^4=-1"

"z=\sqrt[4]{-1}"


The roots of a complex number are also given by a formula "\sqrt[n]{z}=\sqrt[n]{|z|}\left(\cos\frac{\varphi+2\pi k}{n}+i\sin\frac{\varphi+2\pi k}{n}\right)", where "k=\overline{0,n-1}".


"|-1|=1", "\varphi=\pi"


"k=0":

"z_1=\sqrt[4]{-1}=\sqrt[4]{1}\left(\cos\frac{\pi+2\pi\cdot0}{4}+i\sin\frac{\pi+2\pi\cdot0}{4}\right)=1\cdot\left(\cos\frac{\pi}{4}+i\sin\frac{\pi}{4}\right)=\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2}"


"k=1":

"z_2=\sqrt[4]{-1}=\sqrt[4]{1}\left(\cos\frac{\pi+2\pi\cdot1}{4}+i\sin\frac{\pi+2\pi\cdot1}{4}\right)=1\cdot\left(\cos\frac{3\pi}{4}+i\sin\frac{3\pi}{4}\right)=\frac{\sqrt{2}}{2}-i\frac{\sqrt{2}}{2}"


"k=2":

"z_3=\sqrt[4]{-1}=\sqrt[4]{1}\left(\cos\frac{\pi+2\pi\cdot2}{4}+i\sin\frac{\pi+2\pi\cdot2}{4}\right)=1\cdot\left(\cos\frac{5\pi}{4}+i\sin\frac{5\pi}{4}\right)=-\frac{\sqrt{2}}{2}-i\frac{\sqrt{2}}{2}"


"k=3":

"z_4=\sqrt[4]{-1}=\sqrt[4]{1}\left(\cos\frac{\pi+2\pi\cdot3}{4}+i\sin\frac{\pi+2\pi\cdot3}{4}\right)=1\cdot\left(\cos\frac{7\pi}{4}+i\sin\frac{7\pi}{4}\right)=-\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2}"


So "z_1=\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2}", "z_2=\frac{\sqrt{2}}{2}-i\frac{\sqrt{2}}{2}", "z_3=-\frac{\sqrt{2}}{2}-i\frac{\sqrt{2}}{2}", "z_4=-\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2}"